# Differential privacy guarantees of Markov chain Monte Carlo algorithms

### v1

@@ 3.2 Differential privacy of Monte Carlo estimators
We are now concerned with the task of releasing an estimate of an expectation $\pi_{\mathcal{D}}(f):=\int_{\mathrm{E}}f(x)\pi_{\mathcal{D}}(\mathrm{d}x)$ obtained running an MCMC algorithm for $N$ iterations. Specifically, we are given $f:\mathbb{R}^{d}\to\mathbb{R}$ and we simulate a Markov chain $X^{\mathcal{D}}$ with transition kernel $P_{\mathcal{D}}$ to obtain the ergodic average $\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{\mathcal{D}})$. Similarly to the previous section, our strategy to obtain a DP guarantee will be based on an assumption that requires the ergodic average to be close to the truth, i.e. $\pi_{\mathcal{D}}(f)$. Moreover, it will also be necessary to add noise to the ergodic average since for large $N$ it converges to the true expectation of the observable, $\pi_{\mathcal{D}}(f),$ and hence an adversary could distinguish between two adjacent datasets based on the observed ergodic average. To protect against an adversary we then consider the randomised algorithm
$\mathcal{A}({\mathcal{D}})=\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{\mathcal{D}})+L_{\eta},$ (7)
where $L_{\eta}$ is a random variable that is independent of the chain $X^{\mathcal{D}}$ and should satisfy the following assumption.
@@
For any measurable set $\mathsf{B}$, and any $a,b\in\mathbb{R}$ such that $\lvert a-b\rvert\leqslant\eta$, for $\eta>0$, there exist $(\varepsilon,\delta)$ such that
$\mathbb{P}(a+L_{\eta}\in\mathsf{B})\leqslant e^{\varepsilon}\,\mathbb{P}(b+L_{\eta}\in\mathsf{B})+\delta.$ (8)
This condition essentially requires that adding $L_{\eta}$ makes the release of scalars that are close to each other differentially private. A typical choice is to draw $L_{\eta}$ from the Laplacian distribution with scale parameter $\eta/\varepsilon$, which satisfies Assumption 3.5 with parameters $(\varepsilon,0)$ (Dwork et al., 2006).
In order to take advantage of Assumption 3.5, we shall require that the ergodic averages corresponding to two Markov chains for two adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}$ are $\eta$-close with high probability. Without such property, an adversary would be able to distinguish between two adjacent datasets based on the observed Monte Carlo estimator. We formalise this requirement in the next assumption.
@@
Note that in Assumption 3.6, we do not suppose that ${X^{\mathcal{D}}_{n}}$ and $X^{{\mathcal{D}}^{\prime}}_{n}$ are independent and we can consider any coupling between $\nu_{{\mathcal{D}}}P_{{\mathcal{D}}}^{n}$ and $\nu_{{\mathcal{D}}^{\prime}}P_{{\mathcal{D}}^{\prime}}^{n}$.
In order to obtain a guarantee on the DP of $\mathcal{A}$, we shall assume that $L_{\eta}$ satisfies Assumption 3.5 with $\eta$ as given by Assumption 3.6. Notice that the variance of the noise injected in the output to achieve a fixed level of DP $\varepsilon$ increases with $\eta$. In particular, in the case of Laplacian noise we find
$\textnormal{Var}(\mathcal{A}({\mathcal{D}}))=\textnormal{Var}\left(\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{\mathcal{D}})\right)+\frac{\eta}{\varepsilon_{\eta}}\;.$
@@
We are now ready to state our result, that is obtained applying a more general argument that holds for any algorithm of the type $\mathcal{A}({\mathcal{D}})=g_{\mathcal{D}}+L_{\eta}$, where $g_{\mathcal{D}}$ is a random function that depends on ${\mathcal{D}}$ (see Proposition A.5 in Section A.3).
@@
Suppose Assumption 3.6 holds for some $\eta$ and $\tilde{\delta}$, and that $L_{\eta}$ satisfies Assumption 3.5 with the same $\eta$. Then the randomised mechanism $\mathcal{A}$, defined in (7), is $(\varepsilon,\delta+\tilde{\delta})$-differentially private.
@@
The result follows applying Proposition A.5. ∎
Verifying Assumption 3.6 directly would require ad-hoc arguments. Therefore we now introduce two conditions that are sufficient to ensure that Assumption 3.6 holds. We start with a high-probability, non-asymptotic bound on the convergence of the ergodic average to the true value, $\pi_{\mathcal{D}}(f).$
@@
Let $\{P^{n}_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ be a family of Markov kernels and $\{\nu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ a family of initial distributions and let $N\in\mathbb{N}$. Denote by $({X^{\mathcal{D}}_{n}})_{n=1}^{N}$ a Markov chain with transition kernel $P_{\mathcal{D}}$ for any ${\mathcal{D}}\in\mathcal{S}$. Assume that there exist $\tilde{\delta}\in(0,1)$ and $C<\infty$ such that for any ${\mathcal{D}}\in\mathcal{S}$
@@
This assumption can be typically verified using some mixing properties of the family of Markov kernels $\{P^{n}_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$, that should be uniform over ${\mathcal{D}}\in\mathcal{S}$ as stated in Assumption 3.1; see e.g. Paulin (2015); Durmus et al. (2023) and the reference therein.
@@
Suppose Assumptions 3.8 and 3.9 hold for some $C,\tilde{\delta},\gamma_{f}$. Suppose also that $L_{\eta}$ satisfies Assumption 3.5 for $\eta=2C+\gamma_{f}$. Then the randomised mechanism $\mathcal{A}$ defined in (7) is $(\varepsilon,\delta+2\tilde{\delta}-\tilde{\delta}^{2})$-differentially private.
@@
The proof is based on showing that Assumption 3.6 holds under our assumptions. The details can be found in Section A.4. ∎
@@ 4 Characterising the Privacy of Diffusions
In this section we describe our novel proof strategy to obtain DP guarantees of MCMC algorithms.
@@ 4.1 Differential Privacy with Radon-Nikodym Derivatives
We shall start describing our approach considering an abstract randomised algorithm $\mathcal{A}({\mathcal{D}}):\Omega\to\mathcal{X}$ on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, that should be interpreted as the output of an MCMC algorithm such as its $n$-th state or its full path up to the the $n$-th state. The distribution of $\mathcal{A}({\mathcal{D}})$ for any ${\mathcal{D}}\in\mathcal{S}$, denoted as $P_{\mathcal{A}({\mathcal{D}})}$, is defined for any measurable set $\mathsf{B}\subset\mathcal{X}$ as
$P_{\mathcal{A}({\mathcal{D}})}^{\mathbb{P}}(\mathsf{B})=\mathbb{P}(\mathcal{A}({\mathcal{D}})\in\mathsf{B})\;.$
Our strategy to obtain DP of the mechanism $\mathcal{A}$ is based on a change of measure argument. In particular, for every pair of adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ we shall find a probability measure $\mathbb{Q}$ on $(\Omega,\mathcal{F})$ such that the law of $\mathcal{A}({\mathcal{D}})$ under $\mathbb{Q}$ is equal to the law of $\mathcal{A}({\mathcal{D}}^{\prime})$ under $\mathbb{P}$. This is to say,
@@
Now that we have set the framework, we can show how to obtain DP of $\mathcal{A}$ bounding the RN derivative with high probability.
@@
In the next section, we will rely on Lemma 4.1 to obtain DP guarantees for diffusion-based MCMC algorithms, where the RN derivative can be obtained by Girsanov’s theorem. The following lemma is essentially a version of the data processing inequality.
@@
For all results in this paper, bounds in $(\varepsilon,\delta)$-DP can be achieved by converting a Rényi-DP bound, as mentioned in Section 2. However, we presented Lemmas 4.1 and 4.2 separately since Lemma 4.2 applies more generally, and in particular does not require control of the tails of d/$\mathrm{d}\mathbb{P}/\mathrm{d}\mathbb{Q}$.
@@ 4.2 Differential privacy of stochastic differential equations
with initial condition $X^{\mathcal{D}}_{0}=x_{0}\in\mathbb{R}^{d}.$ Here $W_{t}$ is Brownian motion, $f_{\mathcal{D}}:\mathbb{R}^{d}\times\mathcal{X}\to\mathbb{R}^{d}$ is a measurable function, $\kappa:[0,\infty)\to[0,\infty)$ is a function satisfying $\kappa(t)\leqslant t$, $(\eta_{t})_{t\geqslant 0}$ is a stochastic process independent of $(W_{t})_{t\geqslant 0}$ taking values in some space $\mathcal{X}$, and $\beta>0$ is an inverse temperature parameter that scales the noise. The function $\kappa$ allows us to consider discrete time approximations obtained e.g. applying the Euler scheme to a continuous time SDE. For instance, the choice $\kappa(t):=\gamma\lfloor t/\gamma\rfloor$ corresponds to the backwards projection onto the grid $\{n\gamma\}_{n\in\mathbb{N}}$. In this case one has that $X^{\mathcal{D}}_{n\gamma}$ is equal in law to the Markov chain $(x^{\mathcal{D}}_{n})_{n\in\mathbb{N}}$
@@
We assume that (16) has a unique strong solution. This holds when $\kappa$ is a projection onto a grid, and also when $\kappa(t)=t$ under weak conditions (Zhang, 2005).
Now that we have set the framework, we are ready to give our result on the privacy of the entire trajectory of $X^{\mathcal{D}}$ up to a time $T>0$.
Proposition 4.5 (Privacy of the path).
Consider the family of processes (16). Suppose there exists $c>0$ such that for every adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ and $x\in\mathbb{R}^{d},\eta\in\mathcal{X}$ one has
and $C_{1}(T)=c^{2}\beta T$.
Proof.
Fixing adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$, we use Girsanov’s theorem to define a measure $\mathbb{Q}$ under which $(X^{{\mathcal{D}}^{\prime}}_{t})_{t\in[0,T]}$ is equal in distribution to $(X^{{\mathcal{D}}}_{t})_{t\in[0,T]}$ under $\mathbb{P}$. Using bounds on the assumed discrepancy between the drifts (18), we may control the RN derivative d/$\mathrm{d}\mathbb{P}/\mathrm{d}\mathbb{Q}$ in such a way as to obtain the result applying Lemmas 4.1 and 4.2. Full details are presented in Section B.5. ∎
@@
Let $T>0$ and assume there exist constants $L,C,c>0$ such that for every adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ one has
@@
and $(X_{t}^{\mathcal{D}})_{t\in\left[0,T\right]}$ and $(X_{t}^{{\mathcal{D}}^{\prime}})_{t\in\left[0,T\right]}$ are solutions of (16) with drifts $f_{{\mathcal{D}}}$ and $f_{{\mathcal{D}}^{\prime}}$ respectively, driven by the same Brownian motion $(W_{t}){{}_{t}\geqslant 0}$, such that almost surely it holds that
@@
In particular, $Z_{t}$ is given by perturbing the dynamics of $X^{{\mathcal{D}}}_{T}$ on $t\in[T-1,T]$ in such a way that it approaches $X^{{\mathcal{D}}^{\prime}}_{T}$. By the assumption that $X_{t}^{\mathcal{D}}$ and $X_{t}^{{\mathcal{D}}^{\prime}}$ are almost surely close, one may define this perturbation in such a way that it is almost surely bounded. Therefore via Girsanov’s theorem we can find a measure $\mathbb{Q}$ satisfying the assumptions of Lemmas 4.1 and 4.2. Full details are presented in Section B.6. ∎

### v2

@@ 3.2 Differential privacy of Monte Carlo estimators
We are now concerned with the task of releasing an estimate of an expectation $\pi_{\mathcal{D}}(f):=\int_{\mathrm{E}}f(x)\pi_{\mathcal{D}}(\mathrm{d}x)$ obtained running an MCMC algorithm for $N$ iterations. Specifically, we are given an observable $f:\mathrm{E}\to\mathbb{R}$ and we simulate a Markov chain $X^{\mathcal{D}}$ with transition kernel $P_{\mathcal{D}}$ to obtain the ergodic average $\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{\mathcal{D}})$. Similarly to the previous section, our strategy to obtain a DP guarantee will be based on an assumption that requires the ergodic average to be close to the truth, i.e. $\pi_{\mathcal{D}}(f)$.
We consider the randomised algorithm
$\mathcal{A}({\mathcal{D}})=\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{\mathcal{D}})+L,$ (9)
where $L$ is a random variable that is independent of the chain $X^{\mathcal{D}}$ and $f$ is the observable of interest. Here, it is crucial to add noise to the ergodic average to prevent an adversary from distinguishing between two adjacent datasets based on the observed output. In particular, $L$ should satisfy the following assumption, which makes the release of scalars that are close to each other differentially private.
@@
Let $\eta\in(0,\infty)$. There exist $(\varepsilon,\delta)\in\mathbb{R}_{+}\times[0,1)$ such that, for any measurable set $\mathsf{B}$ and any $a,b\in\mathbb{R}$ for which $\lvert a-b\rvert\leqslant\eta$, it holds
$\mathbb{P}(a+L\in\mathsf{B})\leqslant e^{\varepsilon}\,\mathbb{P}(b+L\in\mathsf{B})+\delta.$ (10)
A typical choice is to draw $L$ from the Laplacian distribution with scale parameter $\eta/\varepsilon$, which for any $\eta\in(0,\infty)$ satisfies Assumption˜3.5 with parameters $(\varepsilon,0)$ (Dwork et al., 2006).
In order to take advantage of Assumption˜3.5, we shall require that the ergodic averages for the observable $f$ corresponding to two Markov chains for two adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}$ are $\eta$-close with high probability. Without such a property, an adversary would be able to distinguish between two adjacent datasets based on the observed Monte Carlo estimator. We formalise this requirement in the next assumption.
@@
Note that in Assumption˜3.6, we do not suppose that $X^{\mathcal{D}}_{n}$} and $X^{{\mathcal{D}}^{\prime}}_{n}$ are independent and we can consider any coupling between $\nu_{{\mathcal{D}}}P_{{\mathcal{D}}}^{n}$ and $\nu_{{\mathcal{D}}^{\prime}}P_{{\mathcal{D}}^{\prime}}^{n}$. Moreover, we remark that Assumption˜3.6 should hold only for the observable of interest.
In order to obtain a guarantee on the DP of $\mathcal{A}$, we shall assume that $L$ satisfies Assumption˜3.5 with $\eta$ as given by Assumption˜3.6. Notice that the variance of the noise injected in the output to achieve a fixed level of DP $\varepsilon$ increases with $\eta$. In particular, in the case of Laplacian noise we find
$\textnormal{Var}(\mathcal{A}({\mathcal{D}}))=\textnormal{Var}\left(\frac{1}{N}\sum_{n=1}^{N}f(X_{n}^{\mathcal{D}})\right)+\frac{\eta}{\varepsilon}\;.$
@@
We are now ready to state our result, that is obtained applying a more general argument that holds for any algorithm of the type $\mathcal{A}({\mathcal{D}})=g_{\mathcal{D}}+L$, where $g_{\mathcal{D}}$ is a random function that depends on $\mathcal{D}$} (see Proposition˜A.6 in Section˜A.4).
@@
[…]
@@
The result follows applying Proposition˜A.6. ∎
Verifying Assumption˜3.6 directly would require ad-hoc arguments. Therefore we now introduce two conditions that are sufficient to ensure that Assumption˜3.6 holds. We start with a high-probability, non-asymptotic bound on the convergence of the ergodic average to the true value, $\pi_{\mathcal{D}}(f).$
@@
Let $f:\mathrm{E}\to\mathbb{R}$. Let $\{P^{n}_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ be a family of Markov kernels and $\{\nu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ a family of initial distributions and let $N\in\mathbb{N}$. Denote by $({X^{\mathcal{D}}_{n}})_{n=1}^{N}$ a Markov chain with transition kernel $P_{\mathcal{D}}$ for any ${\mathcal{D}}\in\mathcal{S}$. There exist $\tilde{\delta}\in(0,1)$ and $C<\infty$ such that for any ${\mathcal{D}}\in\mathcal{S}$
@@
This assumption can be typically verified using some mixing properties of the family of Markov kernels $\{P^{n}_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$, that should be uniform over ${\mathcal{D}}\in\mathcal{S}$ as stated in Assumption˜3.1; see e.g. Paulin (2015); Durmus et al. (2023) and the reference therein.
@@
Suppose Assumptions˜3.8 and 3.9 hold for some $C,\tilde{\delta},\gamma_{f}$. Suppose also that $L$ satisfies Assumption˜3.5 for $\eta=2C+\gamma_{f}$. Then the randomised mechanism $\mathcal{A}$ defined in (9) is $(\varepsilon,\delta+2\tilde{\delta}-\tilde{\delta}^{2})$-differentially private.
@@
The proof is based on showing that Assumption˜3.6 holds under our assumptions. The details can be found in Section˜A.5. ∎
@@ 4 Characterising the Privacy of Diffusions
In this section we describe our proof strategy to obtain DP guarantees of MCMC algorithms based on diffusions. We focus here on Radon-Nikodym derivatives, which is essentially an abstraction of the well-known approach using the density function of the random mechanism, see for example in Section 3.1 of Desfontaines & Pejó (2022) or Lemma 7.1.5 of Vadhan (2017) for Lemma 4.1. This framework allows us to calculate privacy parameters via Girsanov’s theorem, which describes a change of measure on the probability space.
@@ 4.1 Differential Privacy with Radon-Nikodym Derivatives
We shall start describing our approach considering an abstract randomised algorithm $\mathcal{A}({\mathcal{D}}):\Omega\to\mathcal{X}$ taking values in some (measurable) space $\mathcal{X}$ and defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. This shall be interpreted in the sequel as the output of an MCMC algorithm such as its $n$-th state or its full path up to the the $n$-th state. Here we stress the dependence on $\mathbb{P}$ and denote the distribution of $\mathcal{A}({\mathcal{D}})$ for any ${\mathcal{D}}\in\mathcal{S}$ as
$P_{\mathcal{A}({\mathcal{D}})}^{\mathbb{P}}(\mathsf{B})=\mathbb{P}(\mathcal{A}({\mathcal{D}})\in\mathsf{B})\;,$
for any measurable set $\mathsf{B}\subset\mathcal{X}.$ Our strategy to obtain DP of the mechanism $\mathcal{A}$ is based on a change of measure argument. In particular, for every pair of adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ we shall find a probability measure $\mathbb{Q}$ on $(\Omega,\mathcal{F})$ such that the law of $\mathcal{A}({\mathcal{D}})$ under $\mathbb{Q}$ is equal to the law of $\mathcal{A}({\mathcal{D}}^{\prime})$ under $\mathbb{P}$. This is to say,
@@
Now that we have set the framework, we can show how to obtain DP of $\mathcal{A}$ bounding the Radon-Nikodym derivative with high probability.
@@
In the next section, we will rely on Lemma˜4.1 to obtain DP guarantees for diffusion-based MCMC algorithms, where the Radon-Nikodym derivative can be obtained by Girsanov’s theorem. The following lemma is essentially a corollary of the data processing inequality, see Theorem 9 in van Erven & Harremos (2014).
@@
For all results in this paper, bounds in $(\varepsilon,\delta)$-DP can be achieved by converting a Rényi-DP bound, as mentioned in Section˜2. However, we presented Lemmas˜4.1 and 4.2 separately since Lemma˜4.2 applies more generally, and in particular does not require control of the tails of $\mathrm{d}\mathbb{P}/\mathrm{d}\mathbb{Q}$.
@@ 4.2 Differential privacy of stochastic differential equations
with initial condition $X^{\mathcal{D}}_{0}=x_{0}\in\mathbb{R}^{d}.$ Here $W_{t}$ is Brownian motion, $f_{\mathcal{D}}:\mathbb{R}^{d}\times\mathcal{Y}\to\mathbb{R}^{d}$ is a measurable function, $\kappa:[0,\infty)\to[0,\infty)$ is a function satisfying $\kappa(t)\leqslant t$, $(\eta_{t})_{t\geqslant 0}$ is a stochastic process independent of $(W_{t})_{t\geqslant 0}$ taking values in some space $\mathcal{Y}$, and $\beta>0$ is an inverse temperature parameter that scales the noise. The function $\kappa$ allows us to consider discrete time approximations obtained e.g. applying the Euler scheme to a continuous time SDE. For instance, the choice $\kappa(t):=\gamma\lfloor t/\gamma\rfloor$ corresponds to the backwards projection onto the grid $\{n\gamma\}_{n\in\mathbb{N}}$. In this case one has that $X^{\mathcal{D}}_{n\gamma}$ is equal in law to the Markov chain $(x^{\mathcal{D}}_{n})_{n\in\mathbb{N}}$
@@
We assume without further commentary that (18) has a unique strong solution. This holds when $\kappa$ is a projection onto a grid, and also when $\kappa(t)=t$ under weak conditions (Zhang, 2005).
@@
Let $T>0$ and assume there exist constants $L,C,c>0$ such that for every adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ one has for any $x,y\in\mathbb{R}^{d}$ and $s\in\mathcal{Y}$
@@
and $(X_{t}^{\mathcal{D}})_{t\in\left[0,T\right]}$ and $(X_{t}^{{\mathcal{D}}^{\prime}})_{t\in\left[0,T\right]}$ are solutions of (18) with drifts $f_{{\mathcal{D}}}$ and $f_{{\mathcal{D}}^{\prime}}$ respectively, driven by the same Brownian motion $(W_{t}){{}_{t}\geqslant 0}$, such that almost surely [^1] Since the measures $\mathbb{P}$ and $\mathbb{Q}$ are absolutely continuous, here and elsewhere we don’t distinguish for which measure the event in question is almost sure. it holds that
@@
In particular, $Z_{t}$ is given by perturbing the dynamics of $X^{{\mathcal{D}}}_{T}$ on $t\in[T-1,T]$ in such a way that it approaches $X^{{\mathcal{D}}^{\prime}}_{T}$. By the assumption that $X_{t}^{\mathcal{D}}$ and $X_{t}^{{\mathcal{D}}^{\prime}}$ are almost surely close, one may define this perturbation in such a way that it is almost surely bounded. Therefore via Girsanov’s theorem we can find a measure $\mathbb{Q}$ satisfying the assumptions of Lemmas˜4.1 and 4.2. Full details are presented in Section˜B.6. ∎
Furthermore, we have the following result on the privacy of the entire trajectory of $X^{\mathcal{D}}$ up to a time $T>0$.
Proposition 4.6 (Privacy of the path).
Consider the family of processes (18). Suppose there exists $c>0$ such that for every adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ and $x\in\mathbb{R}^{d},\eta\in\mathcal{Y}$ one has
and $C_{1}(T)=c^{2}\beta T$.
Proof.
Fixing adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$, we use Girsanov’s theorem to define a measure $\mathbb{Q}$ under which $(X^{{\mathcal{D}}^{\prime}}_{t})_{t\in[0,T]}$ is equal in distribution to $(X^{{\mathcal{D}}}_{t})_{t\in[0,T]}$ under $\mathbb{P}$. Using bounds on the assumed discrepancy between the drifts (24), we may control the Radon-Nikodym derivative $\mathrm{d}\mathbb{P}/\mathrm{d}\mathbb{Q}$ in such a way as to obtain the result applying Lemmas˜4.1 and 4.2. Full details are presented in Section˜B.5. ∎